Bible wants to save R330 284.00 with the next four years. She makes equal monthly deposits at the end of each month starting after one month and her savings accumulate interest at 12.5% per annum.compounded monthly. Calculate the size of each of the monthly deposit


1
Expert's answer
2021-05-07T11:40:43-0400

"FV= C*[\frac{(1+i)^n-1)}{i}]"

"330284.00=C*[\frac{(1+\frac{0.125}{12})^{4*12}-1}{\frac{0.125}{12}}]"

C=R5338.49